# LLM-Generated Heuristics for AI Planning: Do We Even Need Domain-Independence Anymore?

### v3

@@ Abstract
Heuristics are a central component of classical AI planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning. This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions—bypassing the need for handcrafted domain knowledge. We present a method that employs LLMs to generate instance-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language. These heuristics are compiled and integrated into standard heuristic search algorithms, such as greedy best-first search. Our approach achieves competitive, and in many cases state-of-the-art, performance across a broad range of established planning benchmarks. Moreover, it enables the solution of problems that are difficult to express in traditional formalisms, including those with complex numeric constraints or custom transition dynamics. We analyze the trade-offs between generality and performance, and argue that LLM-generated heuristics represent a promising direction for integrating machine learning with symbolic reasoning in planning.
@@ Introduction
Deterministic AI planning, at its core, involves finding a sequence of actions that transitions an agent from an initial state to a goal state, given a formal description of the environment’s dynamics. To search effectively through large state spaces, planners rely on heuristics—functions that estimate the distance to the goal and guide the search toward promising regions.
@@
Classical planning has traditionally focused on general, domain- and problem-independent heuristics, which offer broad applicability and avoid the need for task-specific design (Bonet and Geffner 2001). However, the performance of such heuristics can depend on the representation language and the structural characteristics of the domain. In more expressive settings, general heuristics may face limitations or require adaptation to remain effective in different versions of pddl (cf. Fox and Long 2003; Edelkamp and Hoffmann 2004; Gerevini and Long 2005).
@@
In this work, we propose to automatically generate task-specific heuristic functions from formal problem definitions, using LLMs to generate task-specific heuristics, that are then used by a classic AI search algorithm. In this way, we benefit from the systematic and guaranteed search properties of the classic algorithm, while keeping the generation of the heuristics automatic, with no need for expert manual labor for heuristics design.
@@
We represent each planning task using three components implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state. These are provided as input to the LLM, which returns a heuristic function which can be tailored to the specific problem structure. This heuristic is compiled and used to guide search without further interaction with the LLM.
This approach has several advantages. First, it avoids repeated LLM calls during planning, greatly improving efficiency. Second, it enables the use of transparent, verifiable search components: both the LLM-generated heuristic and the underlying task specification are standard Rust code that can be debugged and tested. Third, it allows us to solve planning problems that involve constructs difficult to express in traditional frameworks—such as recursive goals, complex numeric conditions, or custom transition logic.
We show that LLMs can be used to improve general-purpose planners by using them to generate input-specific heuristics, rather than relying on general-purpose heuristics. We evaluate our approach on a set of standard AI planning benchmarks, translated into Rust representations suitable for LLM prompting. In many domains, the resulting heuristics outperform classical domain-independent methods and compete with state-of-the-art numeric planners. Additionally, our method is able to solve novel planning tasks that lack formal encoding in existing planning languages.
@@
This paper revisits AI planning by heuristic search, by replacing task-independent heuristics with LLM-generated task-specific heuristics. Below, we review related work on LLMs and planning. For a more comprehensive overview of LLMs as planning modelers, see the surveys by (Tantakoun, Zhu, and Muise 2025) and by (Aghzal et al. 2025).
@@
(Valmeekam et al. 2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior. While it seems that LLMs cannot yet generate plans efficiently, our work shows that they can be used effectively as a component of a larger planning paradigm. Namely, LLMs can generate heuristics and search components rather than perform planning directly.
A closely related approach is presented by (Katz et al. 2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls. They propose an alternative, Thought of Search (ToS), in which LLMs are used to generate symbolic search components—successor functions and goal tests. These can then be executed efficiently without further model queries. Although their work focuses on small, finite search problems such as the 24 game and Mini Crosswords, they provide compelling evidence that such an approach can drastically reduce computational cost and improve correctness (solved by BFS). The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (Wei et al. 2022), ReAct (Yao et al. 2023b), ReWOO (Xu et al. 2023), Reasoning via Planning (Hao et al. 2023), Tree of Thoughts (Yao et al. 2023a), Graph of Thoughts (Besta et al. 2024), Reflexion (Shinn et al. 2023), Algorithm of Thought (Sel et al. 2024). The ToS approach runs a Breadth-First Search (BFS) using components—successor functions and goal tests generated by LLM. Our approach extends this intuition by employing heuristic search methods to address problems of significantly greater complexity than those solvable through naive BFS—specifically in AI planning, where transition systems often become infinite due to the presence of unbounded numeric fluents. Consequently, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
A recent study in a similar vein to our work is presented in (Corrêa, Pereira, and Seipp 2025). While their approach targets planning tasks specified in the Planning Domain Definition Language (PDDL), ours departs from this framework by operating directly on successor generators and goal-check functions. This shift affords greater representational flexibility, enabling the modeling of tasks that PDDL cannot express effectively. However, abandoning PDDL’s structured representation also limits our ability to exploit domain structure for heuristic derivation.
@@
LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (Liu et al. 2024), as well as other combinatorial optimization problems (Ye et al. 2024). These approaches typically combine LLMs with evolutionary algorithms to evolve heuristics that guide the optimization process. However, due to fundamental differences in problem structure, these methods are not directly applicable to planning tasks without substantial adaptation.
@@
Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (Fox and Long 2003), among such works are (Guan et al. 2023; Liu et al. 2023; Oswald et al. 2024). For instance, (Silver et al. 2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies. (Guan et al. 2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback. A related concept appears in (Zhou et al. 2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems. Their findings support the idea that LLMs can produce verifiable code with minimal manual intervention. To reduce human involvement even further, we propose generating Rust code instead of Python. Since Rust is a compiled language with strong static typing, much of the burden of code verification can be offloaded to the compiler.
Code Generation Our approach relies on LLMs not to solve problems directly, but to generate search guidance in the form of state evaluation heuristic code. (Zhou et al. 2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code. Our work builds on this insight, but targets search guidance functions within a planning framework. The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (Madaan et al. 2023; Zhang et al. 2023; Chen et al. 2024; Muennighoff et al. 2024; Zhong, Wang, and Shang 2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
@@
Solutions to planning problems are typically found via heuristic search algorithms such as Greedy Best-First Search (GBFS) or A* (Hart, Nilsson, and Raphael 1968; Pearl 1984). These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (Pearl 1984). A heuristic function $h:S\to\mathbb{R}_{\geq 0}$ provides such estimates, often derived using domain-independent approximations, such as relaxations of the planning problem. Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(Fox and Long 2003), but recent work also explores learned heuristics (Oswald et al. 2024; Chen, Thiébaux, and Trevizan 2024). Informative heuristics greatly reduce the number of explored states, making them essential for practical planning. In this work, we use an LLM to generate a heuristic function based on the formal problem description written in a general-purpose programming language (Rust). This heuristic is integrated into a GBFS planner to obtain satisficing solutions.
@@
We start with the overview of our approach. The process begins by accepting two types of input: structured representations provided in pddl2.1 (Fox and Long 2003) format and general problem descriptions given informally or in non-standardized formats. The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function (Izhaki, Green, and Shleyfman 2025). For general problem descriptions, we encode them into Rust according to their specifications. Moreover, the initial and the goal states are specified separately in a JSON file. The manual translation is necessary because current LLMs are not sufficiently reliable for this task, and existing tools lack the maturity required to perform it effectively.
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function, along with a prompt designed to produce a heuristic function in Rust suited to the problem. We utilize three prompts, run in succession, for this process. The strategizing prompt [^1] The exact prompt is provided in the supplementary materials. takes as input the domain definition in Rust and outputs a heuristic concept in English. The second prompt, the unrefined heuristic prompt, produces a Rust heuristic reusable across any problem within the domain. Finally, the optional refined heuristic prompt incorporates both the Rust domain file and a JSON file containing specific input, producing a heuristic for the specific problem instance.
The heuristic, successor generator, and goal-testing functions are then integrated into a standard GBFS framework, where the search is guided by the generated heuristic. To ensure correctness, all solutions are verified using appropriate mechanisms based on the input type. For problems specified in pddl2.1, we rely on a standard pddl verifier—val (Howey, Long, and Fox 2004), which cross-checks the solution against the formal specification of the problem. For general problem descriptions, we verify by applying the actions in the plan sequentially, verify the preconditions are met, and that the last state fulfills goal-test. For the visualization of the workflow see Fig. 1. This workflow seamlessly integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains. By automating key steps, we streamline the planning process and enable the handling of a diverse range of problems with minimal human intervention. The only element that still requires human intervention is the specification of the problem.
Prompt Construction
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function. We generate the heuristic using a Chain-of-Thought (Wei et al. 2022) method consisting of three phases (for the full prompts we refer to the Appendix):
Phase 1 (strategizing) [optional]: In the first phase, we give the LLM a Rust domain implementation (successor generator and goal test functions) showing how the domain is simulated and request an English description of a heuristic that could solve it effectively. We further request the model ignore strict theoretical properties and instead will focus on making the heuristic best at facilitating search when combined with GBFS, by approximating monotonous decrease as the given state gets closer to the goal.
Phase 2 (unrefined heuristic): In the second phase, we show precisely the required signature for the heuristic, and request it to implement the english heuristic above into this format. We also provide a list of guidelines to avoid compilation errors, the LLM from requesting further specification, asking us to complete the heuristic, and to deal with other pitfalls. By the end of this phase we have a heuristic which can be slotted into the search.
Phase 3 (refined heuristic) [optional]: In the last phase, we provide the problem instance JSON file, and request a refinement of the heuristic given the details and parameters of the instance. We let the LLM first strategize how it would change the heuristic and only then implement. This allows for more advanced modifications than if it were to immediately start coding autoregressively. We reckon this has four advantages over using the unrefined heuristic:
@@
    Giving the model another iteration allows it to fix any errors in the initial heuristic.
@@
    It tells the model how significant each part of the problem is. Consider a simple instance of Pacman without ghosts. If a heuristic includes a complex subroutine to quantify the risk from the ghosts, now it can remove that subroutine. Furthermore, if the subroutine initialized GhostDistance=float::max and iterated over ghosts to find the nearest ghost, the distance may have remained infinity and caused errors, now cleared up by removing the subroutine.
@@
    At times, a solution state is obvious from the initial state. In these cases showing the model the instance details allows it to rewrite the heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions. An example of this is given in the supplementary material.
We analyze the impact of including or omitting optional phases in the heuristic generation pipeline in the Ablation Study section.
Baseline Domain-Independent Planners FirstCompilation (fc) TimeSlicedRestarts5 (tsr5)
Rust ENHSP-20 NFD MFF GPT Claude GPT Claude
Domain BFS $h^{\text{md}}$ $^{\text{md}} h+h$$^{\text{mrp}}_{\text{+hj}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5
Block Grouping (20) 0 9 15 19 18 0 2 13 10 5 2 11 4 18 17 15 2 17 13
Counters (20) 4 11 13 20 11 12 15 10 9 10 9 12 6 11 11 11 12 12 11
Delivery (20) 1 11 17 10 12 9 20 14 2 10 3 9 3 18 15 17 9 20 6
Drone (20) 3 16 20 18 15 16 16 15 12 12 7 12 13 16 17 15 13 17 15
Expedition (20) 3 5 3 0 6 5 5 2 1 1 3 3 1 2 2 3 4 4 2
Farming (20) 3 20 20 20 20 15 9 20 18 17 9 14 18 20 20 20 16 20 20
FO-Counters (20) 4 7 9 0 15 6 20 7 6 4 5 8 4 7 7 7 7 8 7
FO-Farming (20) 5 20 20 15 20 11 16 20 15 19 18 20 15 20 18 20 20 20 20
FO-Sailing (20) 0 0 0 0 1 16 11 8 3 0 0 4 1 17 10 1 0 13 1
Hydropower (20) 20 20 4 0 20 6 1 17 19 19 18 20 11 20 19 20 20 20 20
Market Trader (20) 0 18 17 0 20 0 0 9 3 8 9 1 5 18 9 20 17 7 16
Pathways (20) 0 1 0 13 2 2 13 1 1 1 1 0 0 1 1 1 1 1 1
Plant Watering (20) 1 1 20 20 19 20 13 8 2 2 2 3 1 14 8 10 4 9 4
Rover (20) 2 4 10 8 6 4 10 4 0 2 1 3 0 5 1 6 2 5 3
Sailing (20) 0 0 0 20 17 10 2 9 4 1 0 1 1 18 15 1 0 3 3
Settlers (20) 0 0 0 5 0 0 6 1 1 0 0 0 0 2 1 1 0 1 0
TPP (20) 2 20 20 4 4 2 4 16 3 6 6 11 6 20 12 11 13 20 12
Zenotravel (20) 4 12 14 16 11 10 0 17 11 7 8 11 4 20 18 14 15 17 11
$\sum$ (360) 52 175 202 188 217 144 163 191 120 124 101 143 93 247 201 193 155 214 165
Twin Prime (20) 4 4 - - - - - 16 16 12 7 12 8 19 18 19 14 16 15
Pacman (20) 1 14 - - - - - 14 4 8 2 14 7 16 12 15 12 16 12
$\sum$ (400) 57 193 - - - - - 221 140 144 110 169 108 282 231 227 181 246 192
Table 1: Coverage results of the baseline, domain-independent planners, and the LLM-generated heuristics. We test six LLMs: GPT-4.1, GPT-4.1-mini, GPT-4o, GPT-4o-mini, Claude Sonnet 3.5 and Claude Haiku 3.5. FirstCompilation (fc) and TimeSlicedRestarts5 (tsr5) denote the first compiled heuristic and the five generated heuristics where each heuristic got 100 sec for the run, respectively. All configuration except BFS and MFF (that uses EHCS) use GBFS.
@@
Figure 2: Per-instance comparison of the runtime of $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ vs. tsr5 GPT-4.1 unrefined. Points below the diagonal favor our approach. On the left we see time spent on solving the instance overall, including the API calls. On the right we display search time only.
@@
The experiments were conducted on a 13th Gen Intel® Core™ i9-13900 processor running at 2.00 GHz, supported by a 64-bit operating system and 32 GB of RAM. To ensure consistency and reliability, each experiment was constrained by a time limit of 10 minutes and a memory usage limit of 8 GB. We test our approach in the context of satisficing numeric planning. We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (Taitler et al. 2024). [^2] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (Gnad et al. 2023). As a proof-of-concept, we also include two domains that cannot be easily expressed using pddl:
@@
We evaluated our approach against the following SOTA methods on the IPC domains. All of the planners below support linear expressions in both conditions and assignment effects. As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (Chen and Thiébaux 2024), running on the Rust implementation of the problems (Izhaki, Green, and Shleyfman 2025). [^3] https://github.com/AlexTuisov/Rust_search
The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm. It is defined as:
@@
i.e., the distance from $\psi(s)$ to the nearest satisfying value of $\psi$. Chen and Thiébaux (2024) proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS with $h^{\text{mrp}}_{\text{+hj}}$ (Scala et al. 2020), $h^{\text{md}}$ and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ (Chen and Thiébaux 2024). We used the modified version released by Chen and Thiébaux (2024). [^4] https://github.com/hstairs/jpddlplus/tree/socs24-width-and-mq
Metric-FF (MFF) (Hoffmann 2003) is used off-the-shelf. Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators. [^5] https://fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
Numeric-FD (NFD) (Aldinger and Nebel 2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (Kuroiwa, Shleyfman, and Beck 2022). [^6] https://github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see Taitler et al. 2024).

### v4

# Successor-Generator Planning with LLM-generated Heuristics

@@ Abstract
Heuristics are a central component of deterministic planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning. This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions – bypassing the need for handcrafted domain knowledge. We present a method that employs LLMs to generate problem-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language. These heuristics are compiled and integrated into standard heuristic search algorithms, such as greedy best-first search. Our approach achieves competitive, and in many cases state-of-the-art, performance across a broad range of established planning benchmarks. Moreover, it enables the solution of problems that are difficult to express in traditional formalisms, including those with complex numeric constraints or custom transition dynamics. We provide an extensive empirical evaluation that characterizes the strengths and limitations of the approach across diverse planning settings, demonstrating its effectiveness.
@@ Introduction
Deterministic AI planning involves finding a sequence of actions that moves an agent from an initial state to a goal state, given a formal description of the environment’s dynamics (ghallab-et-al:2004). Because state spaces are usually exponential or even infinite, effective search strategies require heuristics – functions that estimate the distance to the goal and help prioritize promising states (pearl:1984; bonet2001planning). Heuristic search has therefore become a central paradigm in classical planning, underpinning many of the most successful planners developed over the past decades (hoffman:jair-2003; helmert:jair-2006; scala-et-al:ijcai-2016; aldinger-nebel:socs-2017).
@@
Classical planning has traditionally focused on domain-independent heuristics, which offer broad applicability and avoid the need for task-specific design (bonet2001planning). However, the performance of such heuristics can vary with the representation language and the structural characteristics of the domain. In more expressive settings, general heuristics may encounter challenges and need adaptation to remain effective across different versions of pddl (mcdermott-et-al:tr-1998; fox-long:jair-2003; edelkamp-hoffmann:tr-2004) (e.g., cf. classical vs. numeric settings (hoffman:aimag-2001; hoffman:jair-2003; helmert-domshlak:icaps-2009; kuroiwa-et-al:jair-2022)). For further discussion of current limitations within pddl see edelkamp:icapsws-2003; rintanen:aaai-20215.
@@
In this work, we propose a method for automatically generating problem-specific heuristic functions from formal problem definitions using LLMs. The generated heuristics are then used by a sound search algorithm, such as greedy best-first search (GBFS). This setup preserves the systematic and provably correct behavior of classical search methods, while making heuristic construction fully automatic and eliminating the need for expert-designed heuristics.
@@
We represent each planning task using three components, implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state (russell-norvig:1995; guan-etal:nips-2023; oswald-etal:icaps-2024). We refer to this representation as an Explicit Successor Generator (ESG). These ESG components, together with a prompt, are provided as input to the LLM, which returns a heuristic function (expressed in the same programming language) tailored to the structure of the given problem. The heuristic is then compiled and used to guide search without any further interaction with the LLM.
This approach offers several benefits. It removes the need for repeated LLM calls during planning—a source of significant inefficiency in prior work (tos:nips-2024; valmeekam-etal:corr-2022; kambhampati-etal:icml-2024). It also provides transparent search components: both the LLM-generated heuristic and the task specification are expressed in a general-purpose programming language (Rust in this work), allowing inspection, debugging, and testing. In addition, it accommodates planning tasks that involve constructs difficult to model in traditional frameworks, including recursive goals, intricate numeric conditions, the creation of new variables, or custom transition logic.
We show that LLM-generated heuristic can strengthen general-purpose ESG planning. Our evaluation includes a set of standard numerical planning benchmarks represented in the ESG form (using Rust), as well as new benchmarks that are challenging to encode in pddl. Across many domains, the resulting heuristics achieve strong performance, matching the level of state-of-the-art numeric planners in numerous cases. Furthermore, the method is able to solve planning tasks that currently lack formal encodings in existing planning languages.
@@
This paper revisits AI planning by heuristic search, by replacing domain-independent heuristics with LLM-generated domain-specific ones. Below, we review related work on LLMs and planning. For a more comprehensive overview of LLMs as planning modelers, see the surveys by (tantakoun-etal:corr-2025) and by (aghzal2025surveylargelanguagemodels).
@@
(valmeekam-etal:corr-2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior. While it seems that LLMs cannot yet generate plans efficiently, our work shows that they can be used effectively as a component of a larger planning paradigm. Namely, LLMs can generate heuristics and search components rather than perform planning directly.
A closely related approach is presented by (tos:nips-2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls. They propose an alternative, Thought of Search (ToS), in which LLMs are used to generate symbolic search components—successor functions and goal tests. These can then be executed efficiently without further model queries. Although their work focuses on small, finite search problems such as the 24 game and Mini Crosswords, they provide compelling evidence that such an approach can drastically reduce computational cost and improve correctness (solved by BFS). The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (wei2022chain), ReAct (react:iclr-2023), ReWOO (rewoo:corr-2023), Reasoning via Planning (rap:emnkp-2023), Tree of Thoughts (tot:nips-2023), Graph of Thoughts (got:aaai-2024), Reflexion (reflection:nips-2023), Algorithm of Thought (aot:icml-2024). The ToS approach runs a breadth-first search (BFS) using components—successor functions and goal tests generated by LLM. Our approach extends this intuition by employing heuristic search methods to address problems of significantly greater complexity than those solvable through naive BFS—specifically in AI planning, where transition systems often become infinite due to the presence of unbounded numeric fluents. Thus, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
A recent study in a similar vein to our work is presented in (correa2025classical). While their approach targets planning tasks specified in pddl, ours departs from this framework by operating directly on successor generators and goal-check functions. This shift affords greater representational flexibility, enabling the modeling of tasks that PDDL cannot express effectively. However, abandoning pddl’s structured representation also limits our ability to exploit domain structure for heuristic derivation.
@@
LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (liu2024evolution), as well as other combinatorial optimization problems (ye2024reevo). These approaches typically combine LLMs with evolutionary algorithms to evolve heuristics that guide the optimization process. However, due to fundamental differences in problem structure, these methods are not directly applicable to planning tasks without substantial adaptation.
@@
Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (fox-long:jair-2003), among such works are (guan-etal:nips-2023; liu-etal:corr-2023; oswald-etal:icaps-2024). For instance, (silver-etal:aaai-2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies. (guan-etal:nips-2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback. A related concept appears in (zhou-etal:iclr-2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems. Their findings support the idea that LLMs can produce verifiable code with minimal manual intervention. To reduce human involvement even further, we propose generating Rust code instead of Python. Since Rust is a compiled language with strong static typing, much of the burden of code verification can be offloaded to the compiler.
Code Generation Our approach relies on LLMs not to solve problems directly, but to generate search guidance in the form of state evaluation heuristic code. (zhou-etal:iclr-2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code. Our work builds on this insight, but targets search guidance functions within a planning framework. The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (madaan-etal:nips-2023; zhang-etal:acl-2023; chen-etal:iclr-2024; muennighoff-etal:iclr-2024; zhong-etal:aclf-2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
@@
Solutions to planning problems are typically found via heuristic search algorithms such as GBFS or A* (hart-et-al:ieee-ssc-1968; pearl:1984). These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (pearl:1984). A heuristic function $h:S\to\mathbb{R}_{\geq 0}$ provides such estimates, often derived using domain-independent approximations, such as relaxations of the planning problem. Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(fox-long:jair-2003), but recent work also explores learned heuristics (oswald-etal:icaps-2024; chen-etal:aaai-2024). Informative heuristics greatly reduce the number of explored states, making them essential for practical planning. In this work, we use an LLM to generate a heuristic function based on the formal problem description written in a general-purpose programming language (Rust). This heuristic is integrated into a GBFS planner to obtain satisficing solutions.
@@
We start with the overview of our approach. The process begins by accepting two types of input: structured representations provided in pddl2.1 (fox-long:jair-2003) format and general problem descriptions given informally or in non-standardized formats. The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function. For general problem descriptions, we encode them into Rust according to their specifications. Moreover, the initial and the goal states are specified separately in a JSON file. The manual translation is necessary because current LLMs are not sufficiently reliable for this task, and existing tools lack the maturity required to perform it effectively. The automatic translations to ESG formalism is left for future work.
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function, along with a prompt designed to produce a heuristic function in Rust suited to the problem. The prompt [^1] The exact prompt for both variations is given in the Appendix. has three components:
@@
    Conditioning the model to be a senior engineer in the GPL used (Anam2025PromptEA) and providing the format the resulting heuristic must follow. This component is spread across both the system prompt and the user message [^2] We include guidance to avoid compilation errors, however the guidance could be far better optimized. This optimization, however, is out of scope for this work.
@@
    Requesting that the LLM generate a heuristic.
@@
    Providing the model with the GPL domain implementation (the ESG).
As an additional variation of our method, the user message may also contain a fourth component:
@@
    The JSON representation of the instance to be solved.
We analyze the impact of instance-specificity in the heuristic generation pipeline in the Ablation Study section.
The heuristic, successor generator, and goal-testing functions are then integrated into a standard GBFS framework, where the search is guided by the generated heuristic. To ensure correctness, all solutions are verified using appropriate mechanisms based on the input type. For problems specified in pddl2.1, we rely on a standard pddl verifier—val (howey-etal:ictai-2004), which cross-checks the solution against the formal specification of the problem. For general problem descriptions, we verify by applying the actions in the plan sequentially, verify the preconditions are met, and that the last state fulfills goal-test. For the visualization of the workflow see Fig. 1. This workflow integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains. Note that the only element of the workflow that still requires human intervention is the specification of the problem.
Baseline Domain-Independent Planners FirstCompilation (fc) UntilSuccess (US) SelfPortfolio-10 (SP-10)
Rust ENHSP-20 NFD MFF GPT Claude GPT Claude GPT Claude
Domain BFS $h^{\text{md}}$ $h^{\text{md}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $P(3h||3n)^{\ddagger}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5
@@
[…]
Table 1: Coverage results of the baseline, domain-independent planners, the best portfolio, and the LLM-generated heuristics. We report our results from GPT-4.1, GPT-5.1, and Claude Sonnet 4.5 with configurations FirstCompilation (fc), UntilSuccess (us), and SelfPortfolio-10 (SP-10). All configuration except BFS and MFF (which uses EHCS) use GBFS. Models marked with are reasoning models. For results on all models and configurations see Appendix. The $P(3h||3n)$ portfolio uniformly divides time among three “best” ENHSP heuristics and three “best” novelty heuristics (chen-thiebaux:socs-2024), where the notion of “best” is derived from performance on the same problem set used in our evaluation, rather than being fixed independently of it.
@@
The experiments were conducted on a 13th Gen Intel® Core™ i9-13900 processor running at 2.00 GHz, supported by a 64-bit operating system and 32 GB of RAM. To ensure consistency and reliability, each experiment was constrained by a time limit of 10 minutes and a memory usage limit of 8 GB. We test our approach in the context of satisficing numeric planning. We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (taitler-etal:aimag-2024). [^3] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (gnad-et-al:icaps-2023). As a proof-of-concept, we also include two domains that cannot be easily expressed using pddl:
@@
(a) Total Time (including failed generation + run pairs)
(b) Expanded States
Figure 2: Per-instance comparisons of the Total Time (left) and expanded states (right) between GPT-4.1 with SelfPortfolio-10 and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$. Points below the diagonal favor our approach. On problems both solve they seem to provide similar levels of heuristic guidance. [^5] A portfolio between the two achieves a coverage of 304, however this is a post-hoc optimization and not a fair comparison. We do however consider Domain-Independent+LLM a fruitful direction fo future work.
Figure 3: Per-instance comparison of the Total Time (generation + run) (up) and expanded states (down) between GPT-5.1 with and without setting the reasoning_effort to high. Points below the diagonal favor high reasoning effort. Allowing increased reasoning effort moderately increases heuristic quality, but comes at a heavy expense of time.
We evaluated our approach against the following SOTA methods on the IPC domains. All of the planners below support linear expressions in both conditions and assignment effects. As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (chen-thiebaux:socs-2024), running on a Rust implementation of the problems based on (Domains_Project).
The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm:
@@
i.e., the distance from $\psi(s)$ to the nearest satisfying value of $\psi$. chen-thiebaux:socs-2024 proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS by (scala-et-al:jair-2020), using the modified version released by chen-thiebaux:socs-2024.666github.com/hstairs/jpddlplus/tree/socs24-width-and-mq. Following (chen-thiebaux:socs-2024) we compare against $^{\text{md}}, h+h$$^{\text{mrp}}_{\text{+hj}}$, $h^{\text{add}}$, their novelty variants $^{\text{md}}_{\langle\text{B, QB}\rangle}, h+h$$^{\text{mrp}}_{\text{+hj}~\langle\text{B, QB}\rangle}$, $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, as well as the three portfolio settings, $P(3h)$ (non-novelty only), $P(3n)$ (novelty only), and $P(3h\,\|\,3n)$ (both). We report of them only $h^{\text{md}}$, the best-performing single heuristic $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, and the best portfolio configuration, $P(3h\,\|\,3n)$, while the full comparison is provided in Appendix.
Metric-FF (MFF) (hoffman:jair-2003) is used off-the-shelf. Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators.777fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
Numeric-FD (NFD) (aldinger-nebel:socs-2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (kuroiwa-et-al:icaps-2022).888github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see taitler-etal:aimag-2024).
@@
LLM-generated heuristics are not always compilable and vary in quality and efficiency generation-to-generation. To address this, we employ three fallback strategies:
In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + all API calls).
In UntilSuccess (us), we operate the same as fc, except we continue even if a compilable heuristic failed.